OpenStudy (help_please101):

Which recursive formula generalizes the given sequence? {3, 7, 11, 15, 19, ...} A an = (an- 1) - 4 B an = 2(an- 1) + 1 C an = (an- 1) +4 D an= 4(an-1) - 5

OpenStudy (coacoapuffprincess):

Hmmmm

OpenStudy (518nad):

look at it

OpenStudy (518nad):

a0,a0+4 3,7

OpenStudy (518nad):

the previous term keeps getting added +4 so

OpenStudy (help_please101):

i know its not A but i don't know which one it is. I have been looking at this question for 30 minutes and i have to show my work also.

OpenStudy (518nad):

k lets say a1=3

OpenStudy (518nad):

as u can see a2=a1+4

OpenStudy (518nad):

a3=a2+4

OpenStudy (518nad):

a4=a3+4

OpenStudy (518nad):

a1=3,a2=7,a3=11,a4=15,...

OpenStudy (518nad):

that n-1 is a subscript that might be confusing

OpenStudy (518nad):

/[A_n=A_{n-1}+4/]

OpenStudy (help_please101):

yeah that part is confusing me

OpenStudy (518nad):

its saying the term u want is the term before + 4

OpenStudy (518nad):

the nth term = the (n-1)th term +4

OpenStudy (help_please101):

I'm still really confused, would the answer be C then??

OpenStudy (help_please101):

Is C correct??